given total cost function (TC)= Q2+2Q+2 and total revenue function (TR)= 8Q-3, then calculate the A. profit maximizing level of output? B. maximum profit?


1
Expert's answer
2021-03-14T19:51:03-0400

A) Let's f[nd MC and MR:


"MC=\dfrac{dTC}{dQ}=2Q+2,""MR=\dfrac{dTR}{dQ}=8."

The profit maximizing level of output occurs when MC=MR:


"2Q+2=8,""2Q=6,""Q=3."

B) By the definition, maximum profit can be found as follows:


"\pi=TR-TC,""\pi=8Q-3-(Q^2+2Q+2)=8Q-3-Q^2-2Q-2,""\pi=-Q^2+6Q-5.""\pi_{max}=-(3)^2+6\cdot3-5=\$4."